During a time period of 0.04 s, a force of average strength of 300 N is applied. How much is the change in momentum of the ball in kg m/s. First calculate the impulse and then the change in momentum.


1
Expert's answer
2021-11-09T10:58:29-0500

(a)

"J=F\Delta t=300\ N\times0.04\ s=12\ N \times s."

(b)

"J=F\Delta t=\Delta p=m\Delta v,""\Delta p=F\Delta t=12\ \dfrac{kg\times m}{s}."